assets at the end of the year a. A patent purchased this year from Miller Co. on January 1 for a cash cost of $6,000. When purchased, the patent had an estimated life of 12 years. b. A trademark was registered with the federal government for $4,000. Management estimated that the trademark could be worth as much as $120,000 because it has an indefinite life. c. Computer licensing rights were purchased this year on January 1 for 42,000. The rights are expected to have a six-year useful life to the wary 1. Compute the acquisition cost of each intangible asset. Acquisitlon Cost Patent Trademark Licensing Rights December 31. (Do not round intermediate calculations.) Expenses Patent Trademark Licensing Rights
Answer #1

1) a) The acquisition cost of each intangible asset is shown as follows:-

The acquisition cost of Patent will be equal to cash paid for purchase of patent (i.e. $6,000).

The acquisition cost of trademark will be equal to the amount at which trademark is registered with the federal government (i.e. $4,000).

The acquisition cost of licensing rights is equal to putrchase cost of $42,000.

Asset Acquisition Cost
Patent $6,000
,Trademarks $4,000
Licensing Rights $42,000

b) Amortization Expense for Patent = Acquisition Cost/Useful Life

= $6,000/12 yrs = $500

The trademarks will not be amortized because it has an indefinite life

Amortization Expense for Licensing Rights = Acquisition Cost/Useful Life

= $42,000/6 yrs = $7,000

Total Amortization expense for current year = $500+$7,000 = $7,500

c)   BLUESTONE COMPANY

Income Statement (Partial)

For the year ending December 31

Amortization Expense $7,500

  BLUESTONE COMPANY

Income Statement (Partial)

At December 31

Intangibles:
Patent $6,000
Trademark $4,000
Licensing Rights, $42,000
Less: Accumulated Amortization ($7,500)

2) a) The total cost of machine will include purchase cost, installation costs and renovation costs prior to use but it will not include repairs cost after production began.

Cost of Machine A = $9,300+$950+$750 = $11,000

Cost of Machine B = $38,500+$2,400+$2,000 = $42,900

Cost of Machine C = $22,300+$1,500+$2,500 = $26,300

b) Depreciation Expense for Machine A = (00Cost - Residual Value)/Useful life

= ($11,000 - $1,300)/5 yrs = $1,940

Deprecition Expense for Machine B per hour = (Cost - Residual Value)/Machine hours capacity

= ($42,900-$900)/20,000 = $2.10 per machine hour

Depereciation Exp. for Machine B for current year = 8,000 hrs*$2.10 per machine hour = $16,800

Depreciation rate for Machine C = (1/10 yrs)*2 = 20%

Depreciation Expense for Machine C = Cost*Double Declining Rate

= $26,300*20% = $5,260

Journal Entries (Amounts in $)

Transaction General Journal Debit Credit
1) Depreciation Exp. - Machine A 1,940
Accumulated Depreciation-Machine A 1,940
2) Depreciation Exp. - Machine B 16,800
Accumulated Depreciation-Machine B 16,800
3) Depreciation Exp. - Machine C 5,260
Accumulated Depreciation-Machine C 5,260